OpenStudy (anonymous):

MEDAL REWARDED Describe how the graph of y = x2 can be transformed to the graph of the given equation. y = (x - 2)2 - 15

OpenStudy (anonymous):

@John_ES

OpenStudy (john_es):

It must displaced in the x axis 2 units and 15 units in the y axis.

OpenStudy (john_es):

Displaced to the right 2 units (x axis), and moved down 15 (y axis) 15 units.

OpenStudy (anonymous):

so "shifted" ?

OpenStudy (john_es):

With equations, if \[f(x)=x^2\]Then \[g(x)=f(x-2)-15=(x-2)^2-15\]

OpenStudy (john_es):

Yes, shifted :)

OpenStudy (anonymous):

so shifted right 2 and down 15 ? :)

OpenStudy (john_es):

Yes, perfect. :) (x-a) means shifted a units to the right (x+a) means shifted a units to the left ;)